OpenStudy (anonymous):

Find the x- and y-intercepts and graph the equation plotting the intercepts. Please show all of your work. Submit your graph to the Dropbox. -2x+5y=-15

OpenStudy (karatechopper):

x int -2x+5(0)=-15 -2x=-15 -2x/-2=-15/-2 x=15/2 y int -2x+5y=-15 -2(0)+5y=-15 5y==-15 5y/5=-15/5 y=-3

OpenStudy (anonymous):

^check that

OpenStudy (karatechopper):

hmmm checking..

OpenStudy (anonymous):

x inter:(a,0) y inter:(0,b)

OpenStudy (karatechopper):

where did i mess up?

OpenStudy (karatechopper):

o i gotcha!

OpenStudy (karatechopper):

(15/2,0) (0,-3)

OpenStudy (karatechopper):